OpenStudy (anonymous):

a rectangular container van needs to be filled with identical cubical balikbayan boxes.If the container van's length ,width and height are 16 ft,4 ft,and 6 ft respectively,while each balikbayan box has an edge of 2 ft, what is the maximum number of balikbayan boxes that can be places?

OpenStudy (anonymous):

compute for the vol of the van's container

OpenStudy (anonymous):

volume of van container = \[16 \times 4 \times 6 \] = 384 volume of one small cube = \[2 \times 2 \times2\] =8 No of cubes = \[384\div 8 = 48\] so it is 48 cubes